Sarsh is designing a rectangular garden in her back yard. The length of the garden is to be three feet less than twice its width. if the perimeter of the garden is sixty feet, what are the width and length of her garden
Found 2 solutions by checkley77, nerdybill:
Answer by checkley77(12569)

LENGTH=2W-3
WIDTH=W
2(2W-3)+2W=60
4W-6+2W=60
6W=60+6
6W=66
W=66/6
W=11 FEET.
L=2*11-3
L=22-3
L=19 FEET
PROOF:
2*19+2*11=60
38+22=60
60=60

Answer by nerdybill(6948)

Sarsh is designing a rectangular garden in her back yard. The length of the garden is to be three feet less than twice its width. if the perimeter of the garden is sixty feet, what are the width and length of her garden
.
Let w = width of garden
then from: "The length of the garden is to be three feet less than twice its width." we have
2w-3 = length of garden
.
Perimeter for any rectangle is
2(length + width)
.
The problem also gave us the perimeter as 60 feet, therefore:
60 = 2(w + (2w-3))
60 = 2(w + 2w - 3)
60 = 2(3w - 3)
60 = 6w - 6
66 = 6w
11 feet = w (width of garden)
.
For the length:
2w-3 = 2(11)-3 = 22-3 = 19 feet (length of garden)